OpenStudy (anonymous):

Is my answer correct? Write an exponential function to model this situation: a population of 420 animals decreases at an annual rate of 21%. Then predict the value of the function after 5 years (to the nearest whole number). My answer: f(x) = 420(1.21)^x; 2541 animals

OpenStudy (anonymous):

420(1.21)^5 = 2541

hero (hero):

The function is decreasing so that means that the value in parentheses would have to be between 0 and 1

OpenStudy (anonymous):

I misread the question, I thought it said it was increasing.

OpenStudy (anonymous):

Would it be, f(x) = 420(0.79)x; 129 animals?

OpenStudy (anonymous):

\[P(t)=420\times (.79)^t\] \[P(5)=420\times (.79)^5=129\] rounded

OpenStudy (anonymous):

yes, 129